Dahn left gas station @ 4:30 and headed south. At 6:00 dazy left the same gas station headed north and traveled 10mph faster than Dahn. at 8:30 Dazy called Dahn and found they were 376 miles apart. How fast was each of them traveling?
Answer by mananth(16949)

Dahn left gas station @ 4:30 and headed south. At 6:00 dazy left the same gas station headed north and traveled 10mph faster than Dahn. at 8:30 Dazy called Dahn and found they were 376 miles apart. How fast was each of them traveling?
Dahn speed = x mph
Dazy speed = x+10 mph
..
In 1.5 hours Dahn has traveled 1.5x miles
...
1.5x + 2.5x +2.5(x+20)= 376
4x+2.5x+50=376
6.5x=326
x= 50.1 mph Dahn speed
..
Liza speed = 70.1 mph